The phase difference between two points on a progressive wave is 3𝜋/4


. What will be the


corresponding path difference?

1
Expert's answer
2022-03-16T18:28:36-0400

The phase difference is related with path difference by an equation

"\Delta\phi=\frac{2\pi}{\lambda}\Delta x"

So

"\Delta x=\frac{\Delta \phi}{2\pi}\lambda=\frac{3\pi/4}{2\pi}\lambda=\frac{3\lambda}{8}"